Lisa Sawrey is an 80-year-old retired businesswoman who has been a widow for 10 years. She uses a walker and has poorly fitting dentures. She was recently admitted to the hospital with pneumonia and also has congestive heart failure and diabetes. She routinely takes several medications to control her blood glucose levels, hypertension, and heart function. In addition to these medications, the physician has recently ordered antibiotics to treat the pneumonia. During an initial nutrition screening, Mrs. Sawrey stated that she had been eating very poorly over the past two weeks. She said that she usually weighs about 125 pounds—a fact that was documented in her medical chart from a previous visit. Although she felt she was losing weight, she didn’t know how much weight she may have lost or when she started losing weight. Upon admission to the hospital, Mrs. Sawrey weighed 110 pounds and was 5 feet 3 inches tall. Her serum albumin level was 3.0 grams per deciliter. A physical exam revealed edema, and several other laboratory tests confirmed that she was retaining fluid. As a result of the nutrition screening, Mrs. Sawrey was referred to a nurse for a nutrition assessment.

1. From the brief description provided, which items in Mrs. Sawrey’s medical history, personal and social history, and food and nutrition history might alert the nurse that this patient is at risk of malnutrition?

2. Identify a healthy body weight for Mrs. Sawrey, and calculate her %UBW and %IBW. What do the results reveal? How does the presence of edema influence your evaluation of Mrs. Sawrey’s weight loss?

3. How might fluid retention alter Mrs. Sawrey’s serum protein levels? What physical symptoms may have indicated that she was retaining excess fluid?

4. What tools can be used to estimate Mrs. Sawrey’s usual food intake? What medical, physical, and personal factors are likely to influence her diet?

5. Describe other types of assessment information that may help the nurse determine whether Mrs. Sawrey should be referred to a registered dietitian.
Answer #1

1. The items in Mrs. Sawrey’s medical history, personal and social history, and food and nutrition history might alert the nurse that this patient is at risk of malnutrition are :-

- weight loss

- edema

- albumin 3gm

- diabetes

- poor intake

2. Height of the patient is 5feet 3inch = 160cm

Weight of the patient is 110 lbs = 110×0.45kg = 49.5kg

% IBW = current weight / ideal weight × 100

Ideal body weight = 106 lbs + 6(3) = 106+ 18 = 124 lbs

= 110lbs / 124 × 00

= 110/ 124× 100

= 88.7

% UBW = current body weight / UBW × 100

= 110lbs / 125 lbs × 100

= 110/125 × 100

= 88%

3. Fluid retention or edema occurs due to deficiency of albumin . Hence , protein level decreases due to fluid retention . Physical symptoms that have indicated that she was retaining excess fluid are edema, and several other laboratory tests confirmed that she was retaining fluid.

4. Tools that can be used to estimate Mrs. Sawreys usual food intake are :-

Food record

24hr recall

FFQ

Somw of the factors that can affect her diet are :-

Dentures

Poor intake

Medical conditions like diabetes

No one to take care

Medication